# Korteweg-de Vries Integrals for Modified Black Hole Potentials: Instabilities and other Questions

### v1

@@ 7 Conclusions and Discussion
In Sec. 6, we have adopted the BH moment problem perspective introduced in Refs. [56, 57] (and reviewed in Sec. 3), which allows us to obtain the greybody factors entirely in terms of the KdV integrals, to study the interplay between the stability properties of the KdV integrals and the stability of the greybody factors. This seems to be of particular relevance in the light of some recent works showing how the amplitude of the ringdown signal may be modeled with greybody factors as an alternative to an expansion in QNMs [77, 79]. Our study shows preliminary evidence that the instability of the (higher) KdV integrals only affects the greybody factors locally, while the integrated error we have defined is actually well below the instability zone. In conclusion, there seems to be a very interesting qualitative relation between the KdV integrals and the QNM instability while, at the same time, the moment problem perspective may be the mathematical/analytical tool to explain, starting from the KdV integrals, the stability of the greybody factors. However, further in-depth studies, both analytical and numerical, are required to make stronger statements and establish a solid link between the integrable structures appearing in the perturbative dynamics of BHs and their spectral properties.
@@ Acknowledgments
ML and CFS have been supported by contract PID2019-106515GB-I00/AEI/10.13039/501100011033 (Spanish Ministry of Science and Innovation) and 2017-SGR-1469 (AGAUR, Generalitat de Catalunya). This work was also partially supported by the program Unidad de Excelencia María de Maeztu CEX2020-001058-M (Spanish Ministry of Science and Innovation). ML has also been supported by a Juan de la Cierva contract FJC2021-047289-I funded by program MCIN/AEI/10.13039/501100011033 (Spanish Ministry of Science and Innovation) and by NextGenerationEU/PRTR (European Union). AMA was partially supported by the JAE-ICU program (2023-ICE-07) to support the introduction to scientifc research, linked to the María de Maeztu contract mentioned above He has also been supported by Programme RI-24 - researcher in gravitational-wave astronomy (Research and Innovation Program linked to Component 23, Investment 1, within the framework of the Recovery, Transformation and Resilience Plan, funded by the European Union – Next Generation EU, corresponding to the years 2024 and 2025). We have used the Computer Algebra Systems Mathematica [121] to carry out most of the computations necessary for this paper.
@@ Appendix A Scattering Problem for the Pöschl-Teller Potential
[…]
@@ Appendix C KdV Densities: Reduction of the Total Derivative Order
[…]

### v3

@@ 7 Conclusions and Discussion
In Sec. 6, we have adopted the BH moment problem perspective introduced in Refs. [56, 57] (and reviewed in Sec. 3), which allows us to obtain the greybody factors entirely in terms of the KdV integrals, to study the interplay between the stability properties of the KdV integrals and the stability of the greybody factors. This seems to be of particular relevance in the light of some recent works showing how the amplitude of the ringdown signal may be modeled with greybody factors as an alternative to an expansion in QNMs [77, 78, 79]. Our study shows: i) that the stability/instability of low/high KdV integrals correctly accounts for the stability/instability of the logarithm of the greybody factors (the moment problem distribution) at low/high frequencies; ii) preliminary evidence that the instability of the (higher) KdV integrals only affects the greybody factors locally, while the integrated error we have defined is actually well below the instability zone; iii) that the moment problem perspective to the greybody factors in terms of the KdV integrals seems to account for the high frequency instability of the reflectivity and to relate it to the instability of the higher order KdV integrals. In conclusion, there seems to be a very interesting qualitative relation between the KdV integrals and the QNM instability while, at the same time, the moment problem perspective may be the mathematical/analytical tool to explain, starting from the KdV integrals, the stability of the greybody factors. However, further in-depth studies, both analytical and numerical, are required to make stronger statements and establish a solid link between the integrable structures appearing in the perturbative dynamics of BHs and their spectral properties.
@@ Acknowledgments
ML and CFS have been supported by contract PID2019-106515GB-I00/AEI/10.13039/501100011033 (Spanish Ministry of Science and Innovation) and 2017-SGR-1469 (AGAUR, Generalitat de Catalunya). This work was also partially supported by the program Unidad de Excelencia María de Maeztu CEX2020-001058-M (Spanish Ministry of Science and Innovation). ML has also been supported by a Juan de la Cierva contract FJC2021-047289-I funded by program MCIN/AEI/10.13039/501100011033 (Spanish Ministry of Science and Innovation) and by NextGenerationEU/PRTR (European Union). AMA was supported by the JAE-ICU program (2023-ICE-07) to support the introduction to scientifc research, linked to the María de Maeztu contract mentioned above; by the Programme RI-24 - researcher in gravitational-wave astronomy (Research and Innovation Program linked to Component 23, Investment 1, within the framework of the Recovery, Transformation and Resilience Plan, funded by the European Union – Next Generation EU, corresponding to the years 2024 and 2025); by the Spanish Agencia Estatal de Investigación grants PID2022-138626NB-I00, RED2024-153978-E, RED2024-153735-E, funded by MICIU/AEI/10.13039/501100011033 and the ERDF/EU; and by the Comunitat Autónoma de les Illes Balears through the Conselleria d’Educació i Universitats with funds from the European Union - NextGenerationEU/PRTR-C17.I1 (SINCO2022/6719) and from the European Union - European Regional Development Fund (ERDF) (SINCO2022/18146). We have used the Computer Algebra Systems Mathematica [122] to carry out most of the computations necessary for this paper.
@@ Appendix A Scattering Problem for the Pöschl-Teller Potential
$\psi$ $=$ $A\,2^{\frac{ik}{\alpha}}\left[1-\tanh^{2}(\alpha x)\right]^{-\frac{ik}{2\alpha}}{}_{2}F_{1}\left(\lambda-\frac{ik}{\alpha},1-\lambda-\frac{ik}{\alpha},1-\frac{ik}{\alpha};\frac{1-\tanh(\alpha x)}{2}\right)$ (A.1)
@@ Appendix C KdV Densities: Reduction of the Total Derivative Order
$\mathcal{K}_{n}=\int^{+\infty}_{-\infty}dx\,\kappa_{n}(x)\,,\quad n=1,\ldots\,\,.$ (C.3)